OpenStudy (luigi0210):

Evaluate using a trig sub:

OpenStudy (luigi0210):

\[\LARGE \int \frac{x^3}{\sqrt{1-x^2}}~dx\]

OpenStudy (loser66):

let x = sint dx = cos dt x^2 = sin^2 1-x^2 = cos^2 sqrt (1-x^2) = cos plug all integrand: sin^3 cos dt/cos= sin^3 = (1-cos^2) sin= sin-cos^2sin the second term let u = cos du = -sin it becomes -u^2du I think you understand my short cut

Parth (parthkohli):

whenever given \(\sqrt{1 - x^2}\) in the denom, your first approach should be subbing sin or cos.

OpenStudy (luigi0210):

Thank you guys :)